# Statistical Properties of a Polarization Vector’s Ellipticity Angle

### v1

@@ 3.2 The Q-V Correlation Coefficient
Equation 24 is general in the sense that no assumption has been made regarding the statistical character of the mode intensity fluctuations in its derivation. The mode intensities are assumed to be Gaussian random variables in the analysis that follows.
@@ 3.3 General Joint Probability Density
where $\sigma_{\rho}=\sqrt{1+\rho^{2}}$, and $h(\psi,\chi)$ is
@@
Two examples of Equation 25 are shown in Figure 6. In both cases, the joint density is bimodal due to the fluctuations in the intensities of the OPMs. The top panel of the figure shows $f(\psi,\chi)$ when $s=0$, $\rho=4$, and $\chi_{o}=0^{\circ}$. In this example, each component of the joint density is nearly circularly symmetric. The bottom panel of the figure shows $f(\psi,\chi)$ when $s=1$, $\rho=3$, and $\chi_{o}=30^{\circ}$. The component at $\psi=0^{\circ}$ and $\chi\simeq 30^{\circ}$ is comprised of data samples with $X_{A}>X_{B}$. Similarly, the component at $\psi=\pm 90^{\circ}$ and $\chi\simeq-30^{\circ}$ is comprised of data samples with $X_{A}<X_{B}$. The component at $\psi=0^{\circ}$ is the more prominent of the two because the intensity of mode A is greater than that of mode B on average due to $s=1$. Each component is symmetric in PA and asymmetric in EA. The shapes of the two components become mirror images of one another about $\chi=0^{\circ}$ when $s=0$ and are identical and completely symmetric when $s=0$, $\chi_{o}=0^{\circ}$, and $\rho\gg 1$. The joint density becomes unimodal when the mean amplitude of the polarization vector exceeds the polarization fluctuations (i.e., when $s>\sqrt{1+\rho^{2}}$).
3.4 General Probability Density of the Ellipticity Angle
@@
Figure 7: Examples of the general probability density function (pdf) of the EA for incoherent OPMs. The top panel shows examples of the pdf for different values of the polarization fluctuation ratio, $\rho$, when the intrinsic EA and SNR are held constant at $\chi_{o}=30^{\circ}$ and $s=10$, respectively. The bottom panel shows example pdfs for different values of SNR when the intrinsic EA and the polarization fluctuation ratio are held constant at $\chi_{o}=15^{\circ}$ and $\rho=4$. The dotted vertical lines denote the intrinsic EAs of the OPMs, $\pm\chi_{o}$.
@@
As in Section 2.2, the general pdf of $\chi$ can be found by numerically integrating the joint probability density given by Equation 25 over $\psi$. Examples of this general pdf are shown in Figure 7. The bottom panel of the figure shows the pdf for different values of $s$ when the intrinsic EA of the dominant polarization mode and the polarization fluctuation ratio are held constant at $\chi_{o}=15^{\circ}$ and $\rho=4$, respectively. The pdf is bimodal when the polarization fluctuations exceed the mean polarization ($s<\sqrt{1+\rho^{2}}$) and becomes unimodal as the mean polarization exceeds the fluctuations ($s>\sqrt{1+\rho^{2}}$). The top panel of the figure shows the pdf when the mean polarization is much larger than the fluctuations. These pdfs have different values of $\rho$, but the same values of the intrinsic EA ($\chi_{o}=30^{\circ}$) and polarization SNR ($s=10$). The pdfs are unimodal and almost indistinguishable from one another. (Note that the range of the abscissa in the top panel is half that of the bottom panel). The pdf with $\rho=0$ is nearly symmetric, while the other pdfs have tails that extend toward $\chi=0$. The pdf becomes Gaussian as the mean polarization becomes even larger with respect to the polarization fluctuations. The height of the Gaussian pdf increases and its width decreases with increasing SNR. In this instance, the emission is dominated by one polarization mode.
3.5 Probability Density of the Ellipticity Angle for OPMs with Equal Mean Intensities
@@
Figure 8: Example probability density functions (pdfs) of the EA for incoherent OPMs with the same mean intensity ($s=0$). The top panel shows examples of the pdf for different values of the polarization fluctuation ratio, $\rho$, when the intrinsic EA is held constant at $\chi_{o}=22.5^{\circ}$. The dotted vertical lines denote the intrinsic EAs of the OPMs, $\pm\chi_{o}$. The bottom panel shows example pdfs for different values of $\chi_{o}$ when the polarization fluctuation ratio remains fixed at $\rho=5$.
@@
The function $_{2}F_{1}(a,b;c;z)$ in Equation 28 is the Gauss hypergeometric function with numerical parameters $a$, $b$, and $c$ and argument $z$. This pdf is unimodal with a peak at $\chi=0$. The full width of the EA pdf at half its maximum value (FWHM) approaches $w_{50}=\arcsin(1/\rho)$ as $\rho$ becomes large. An example of this pdf is shown by the dashed blue line in the bottom panel of Figure 8. The pdf of the PA that accompanies $f_{\chi}(\chi)$ in this case is (Equation 17 of McKinnon & Stinebring 1998)
@@
Figure 9: Dependence of the component separation in the EA probability density function (pdf) upon the polarization fluctuation ratio, $\rho$, for different values of the intrinsic EA, $\chi_{o}$. The dotted horizontal lines denote the actual difference in mode EAs, $2\chi_{o}$. In each case, the component separation asymptotically approaches $2\chi_{o}$ as $\rho$ increases.
@@ 4 Comparison with Observations
Both the joint probability density of the polarization angles and the EA pdf derived in Section 3 can be compared to the results of single pulse polarization observations of pulsars to determine if they are reasonable representations of what is actually observed.
4.1 Observed Probability Density of the Ellipticity Angle
@@
Figure 10: Comparison of theoretical EA pdfs with histograms of the EA observed in PSR B2020+28 at 1404 MHz. The top panel shows the EA histogram recorded near the center of the pulse at pulse phase bin (ppb) 74 in Figure 2a of McKinnon (2004). The continuous red line represents the result of a chi-squared fit of the histogram to the theoretical pdf derived numerically from Equation 25. The parameters of the fit are annotated in the panel. The bottom panel compares the EA histogram recorded at the trailing peak of the pulse (ppb 100) with its best fit pdf.
@@ 4.2 Observed Joint Probability Density of the Polarization Angles
Dyks et al. (2021) conducted single pulse polarization observations of PSR B1451-68 with the Parkes radio telescope at a frequency of 1369 MHz. Their Figure 1 shows OPMs occurring across the entirety of the pulsar’s pulse. The mode EAs also vary systematically between $\chi\simeq-\pi/4$ and $\chi\simeq\pi/4$, tracing a bow-tie shape across the pulse. The figure also reveals polarization properties that are consistent with the OPM statistical model summarized in Section 3. Specifically, the instantaneous polarization fraction of the emission can be high, while the polarization fraction on average is low. Their Figure 2 shows color-coded Hammer projections of the PAs and EAs recorded at specific pulse longitudes on the Poincaré sphere. The projections are generally bimodal, suggesting that the polarization fluctuations exceed the mean polarization almost everywhere within the pulse.
@@
Figure 11 shows simulations of the PA-EA joint density at two locations within the pulsar’s pulse. The OPM statistical model was used in the simulations. Each row of the figure is a Lambert equal area projection of the two hemispheres of the Poincaré sphere for a single simulation. The simulations were generated by estimating constant values of $\chi_{o}$, $s$, and $\rho$ from the observations and creating 10,000 samples each of the Stokes parameters $Q$, $U$, and $V$ from Equations 18-20. The simulation modeled the mode intensities as Gaussian random variables. The intrinsic value of the dominant mode PA was set to $\psi_{o}=0$ to simplify the data presentation. Individual samples of the PA and EA were calculated from sets of $Q$, $U$, and $V$ and projected on the Poincaré sphere. [^1] Note that the Poincaré sphere depicts the azimuth, $\phi=2\psi$, and latitude, $\lambda=2\chi$, of the polarization vector. The view of the sphere on the left half of the projection is along the +Q+Q axis of the sphere, and the view on the right half is along the Q-Q axis.
@@
Figure 11: Simulations of the PA-EA joint probability density observed in PSR B1451-68. Each row shows the Lambert equal area projection of the two hemispheres of the Poincaré sphere for a single simulation. The top row shows the results from a simulation of the data recorded at a pulse longitude of $178.1^{\circ}$ (see Figure 2 of Dyks et al. 2021). The parameters used in the simulation were $\chi_{o}=2^{\circ}$, $s=1.8$, and $\rho=6$. The bottom row shows the results from a simulation of the data recorded at a longitude of $188.7^{\circ}$. The parameters used in this simulation were $\chi_{o}=-13^{\circ}$, $s=1$, and $\rho=3.8$. The contour levels shown in both rows are drawn at 0.2, 0.4, 0.6, and 0.8 of the joint density’s peak value.
@@
The top row of Figure 11 is a simulation of the PA-EA joint density at a pulse longitude of $178.1^{\circ}$ in PSR B1451-68. This simulation should be compared to the right center panel in Figure 2 of Dyks et al. (2021). The observed joint density consists of two, circularly-symmetric components. They reside at near-antipodal points close to the equatorial plane of the Poincaré sphere, indicating that the polarization modes are approximately orthogonal and predominantly linearly polarized. One component is more prominent than the other. The model parameters estimated from these observed properties and used in the simulation were $\chi_{o}=2^{\circ}$, $s=1.8$, and $\rho=6$. The results of the simulation are qualitatively consistent with what is observed. The simulation also resembles the PA-EA joint density observed in the trailing outrider of PSR B0329+54 (see the bottom right corner of Figure 2 in Edwards & Stappers 2004).
@@ 5 Comparison of the Statistical Properties of the Polarization Angles
A comparison of the statistical properties of the EA derived in Sections 2 and 3 with those of the PA underscores their similarities and differences. When the amplitude of the polarization vector is constant, the pdf of the PA is always symmetric about the intrinsic PA, $\psi_{o}$. Consequently, the mean of the pdf is always equal to $\psi_{o}$. The mean PA is not biased by the instrumental noise, provided the noise in the Stokes parameters $Q$ and $U$ is independent and equal in magnitude (Montier et al., 2015a, 2015b). The standard deviation of the pdf, $\sigma_{\psi}$, is independent of $\psi_{o}$ and varies inversely with the SNR in linear polarization. The pdf becomes Gaussian as the SNR becomes large, at which point the confidence limits of the pdf are well approximated by $\sigma_{\psi}$. In contrast, the pdf of the EA is generally asymmetric. The standard deviation of the pdf can depend upon the value of the intrinsic EA, $\chi_{o}$, and varies inversely with the SNR in total polarization. The measured EA is biased by the instrumental noise, particularly at low SNR and large values of $\lvert\chi_{o}\rvert$. Since the pdf is asymmetric, its confidence limits are also asymmetric about the mean EA. As the SNR increases, the pdf becomes Gaussian when the SNR and intrinsic EA satisfy the condition $s>3/\cos(2\chi_{o})$, at which point the mean EA is equal to $\chi_{o}$, $\sigma_{\chi}$ is independent of $\chi_{o}$, and the confidence limits of the pdf are well approximated by $\sigma_{\chi}$. The pdf is never symmetric when $\chi_{o}=\pi/4$.
The pdf of the PA is generally bimodal when the amplitude of the polarization vector fluctuates due to the superposition of incoherent OPMs. The two components of the pdf are separated precisely by $\Delta\psi=\pi/2$. Both components are symmetric about their respective centroids with widths scaling inversely as the ratio of the polarization fluctuations relative to the instrumental noise, $\rho$. The pdf of the EA is also bimodal. The two components of the pdf are generally asymmetric about their respective peaks, and their separation tends to be less than $2\chi_{o}$, particularly at large values of $\lvert\chi_{o}\rvert$. As $\rho$ increases, the width of the components decreases, and their separation approaches $2\chi_{o}$. The pdf becomes unimodal when the polarization fluctuations are small relative to either the mean polarization or the instrumental noise.
@@ 6 SUMMARY
The Stokes parameters are covariant when the radio emission from pulsars is comprised of superposed, incoherent OPMs. In this instance, the joint probability density of the polarization angles and the individual pdfs of the EA and PA are unimodal if the mean polarization exceeds the polarization fluctuations and are bimodal if the fluctuations exceed the mean. When the mean intensities of the modes are equal, the widths of the two components in the EA’s bimodal pdf vary inversely with the polarization fluctuations relative to the instrumental noise. The component separation approaches twice the intrinsic EA as the fluctuations increase.
Appendix A Confidence Limits of the Ellipticity Angle

### v2

@@ 3.2 The Q-V Correlation Coefficient
Equation 24 is general, in the sense that no assumption has been made regarding the statistical character of the mode intensity fluctuations in its derivation. The mode intensities are assumed to be Gaussian random variables in the analysis that follows.
@@ 3.3 General Joint Probability Density
where $\sigma_{\rho}=\sqrt{1+\rho^{2}}$ and $h(\psi,\chi)$ is
@@
Two examples of Equation 25 are shown in Figure 6. In both cases, the joint density is bimodal, due to the fluctuations in the intensities of the OPMs. The top panel of the figure shows $f(\psi,\chi)$ when $s=0$, $\rho=4$, and $\chi_{o}=0^{\circ}$. In this example, each component of the joint density is nearly circularly symmetric. The bottom panel of the figure shows $f(\psi,\chi)$ when $s=1$, $\rho=3$, and $\chi_{o}=30^{\circ}$. The component at $\psi=0^{\circ}$ and $\chi\simeq 30^{\circ}$ is comprised of data samples with $X_{A}>X_{B}$. Similarly, the component at $\psi=\pm 90^{\circ}$ and $\chi\simeq-30^{\circ}$ is comprised of data samples with $X_{A}<X_{B}$. The component at $\psi=0^{\circ}$ is the more prominent of the two, because the intensity of mode A is greater than that of mode B on average, due to $s=1$. Each component is symmetric in PA and asymmetric in EA. The shapes of the two components become mirror images of one another about $\chi=0^{\circ}$ when $s=0$ and are identical and completely symmetric when $s=0$, $\chi_{o}=0^{\circ}$, and $\rho\gg 1$. The joint density becomes unimodal when the mean amplitude of the polarization vector exceeds the polarization fluctuations (i.e., when $s>\sqrt{1+\rho^{2}}$).
3.4 General Probability Density of the EA
@@
Figure 7: Examples of the general pdf of the EA for incoherent OPMs. The top panel shows examples of the pdf for different values of the polarization fluctuation ratio, $\rho$, when the intrinsic EA and SNR are held constant at $\chi_{o}=30^{\circ}$ and $s=10$, respectively. The bottom panel shows example pdfs for different values of SNR when the intrinsic EA and the polarization fluctuation ratio are held constant at $\chi_{o}=15^{\circ}$ and $\rho=4$. The dotted vertical lines denote the intrinsic EAs of the OPMs, $\pm\chi_{o}$.
@@
As in Section 2.2, the general pdf of $\chi$ can be found by numerically integrating the joint probability density given by Equation 25 over $\psi$. Examples of this general pdf are shown in Figure 7. The bottom panel of the figure shows the pdf for different values of $s$ when the intrinsic EA of the dominant polarization mode and the polarization fluctuation ratio are held constant at $\chi_{o}=15^{\circ}$ and $\rho=4$, respectively. The pdf is bimodal when the polarization fluctuations exceed the mean polarization ($s<\sqrt{1+\rho^{2}}$) and becomes unimodal as the mean polarization exceeds the fluctuations ($s>\sqrt{1+\rho^{2}}$). The top panel of the figure shows the pdf when the mean polarization is much larger than the fluctuations. These pdfs have different values of $\rho$, but the same values of the intrinsic EA ($\chi_{o}=30^{\circ}$) and polarization SNR ($s=10$). The pdfs are unimodal and almost indistinguishable from one another. (Note that the range of the abscissa in the top panel is half that of the bottom panel.) The pdf with $\rho=0$ is nearly symmetric, while the other pdfs have tails that extend toward $\chi=0$. The pdf becomes Gaussian as the mean polarization becomes even larger with respect to the polarization fluctuations. The height of the Gaussian pdf increases and its width decreases with increasing SNR. In this instance, the emission is dominated by one polarization mode.
3.5 Probability Density of the EA for OPMs with Equal Mean Intensities
@@
Figure 8: Example pdfs of the EA for incoherent OPMs with the same mean intensity ($s=0$). The top panel shows examples of the pdf for different values of the polarization fluctuation ratio, $\rho$, when the intrinsic EA is held constant at $\chi_{o}=22.5^{\circ}$. The dotted vertical lines denote the intrinsic EAs of the OPMs, $\pm\chi_{o}$. The bottom panel shows example pdfs for different values of $\chi_{o}$ when the polarization fluctuation ratio remains fixed at $\rho=5$.
@@
The function $_{2}F_{1}(a,b;c;z)$ in Equation 28 is the Gauss hypergeometric function, with numerical parameters $a$, $b$, and $c$ and argument $z$. This pdf is unimodal, with a peak at $\chi=0$. The full width of the EA pdf at half its maximum value (FWHM) approaches $w_{50}=\arcsin(1/\rho)$ as $\rho$ becomes large. An example of this pdf is shown by the dashed blue line in the bottom panel of Figure 8. The pdf of the PA that accompanies $f_{\chi}(\chi)$ in this case is (Equation 17 of McKinnon & Stinebring 1998)
@@
Figure 9: Dependence of the component separation in the EA pdf upon the polarization fluctuation ratio, $\rho$, for different values of the intrinsic EA, $\chi_{o}$. The dotted horizontal lines denote the actual difference in mode EAs, $2\chi_{o}$. In each case, the component separation asymptotically approaches $2\chi_{o}$ as $\rho$ increases.
@@ 4 Comparison with Observations
Both the joint probability density of the polarization angles and the EA pdf derived in Section 3 can be compared to the results of single-pulse polarization observations of pulsars to determine if they are reasonable representations of what is actually observed.
4.1 Observed Probability Density of the EA
@@
Figure 10: Comparison of theoretical EA pdfs with histograms of the EA observed in PSR B2020+28 at 1404 MHz. The top panel shows the EA histogram recorded near the center of the pulse at pulse phase bin (ppb) 74 in Figure 2a of McKinnon (2004). The continuous red line represents the result of a chi-squared fit of the histogram to the theoretical pdf derived numerically from Equation 25. The parameters of the fit are annotated in the panel. The bottom panel compares the EA histogram recorded at the trailing peak of the pulse (ppb 100) with its best-fit pdf.
@@ 4.2 Observed Joint Probability Density of the Polarization Angles
Dyks et al. (2021) conducted single-pulse polarization observations of PSR B1451-68 with the Parkes radio telescope at a frequency of 1369 MHz. Their Figure 1 shows OPMs occurring across the entirety of the pulsar’s pulse. The mode EAs also vary systematically between $\chi\simeq-\pi/4$ and $\chi\simeq\pi/4$, tracing a bow tie shape across the pulse. The figure also reveals polarization properties that are consistent with the OPM statistical model summarized in Section 3. Specifically, the instantaneous polarization fraction of the emission can be high, while the polarization fraction on average is low. Their Figure 2 shows color-coded Hammer projections of the PAs and EAs recorded at specific pulse longitudes on the Poincaré sphere. The projections are generally bimodal, suggesting that the polarization fluctuations exceed the mean polarization almost everywhere within the pulse.
@@
Figure 11 shows simulations of the PA-EA joint density at two locations within the pulsar’s pulse. The OPM statistical model was used in the simulations. Each row of the figure is a Lambert equal-area projection of the two hemispheres of the Poincaré sphere for a single simulation. The simulations were generated by estimating constant values of $\chi_{o}$, $s$, and $\rho$ from the observations and creating 10,000 samples each of the Stokes parameters $Q$, $U$, and $V$ from Equations 18-20. The simulation modeled the mode intensities as Gaussian random variables. The intrinsic value of the dominant mode PA was set to $\psi_{o}=0$ to simplify the data presentation. Individual samples of the PA and EA were calculated from sets of $Q$, $U$, and $V$ and projected on the Poincaré sphere. [^1] Note that the Poincaré sphere depicts the azimuth, $\phi=2\psi$, and latitude, $\lambda=2\chi$, of the polarization vector. The view of the sphere on the left half of the projection is along the +Q+Q axis of the sphere, and the view on the right half is along the Q-Q axis.
@@
Figure 11: Simulations of the PA-EA joint probability density observed in PSR B1451-68. Each row shows the Lambert equal-area projection of the two hemispheres of the Poincaré sphere for a single simulation. The top row shows the results from a simulation of the data recorded at a pulse longitude of $178.1^{\circ}$ (see Figure 2 of Dyks et al. 2021). The parameters used in the simulation were $\chi_{o}=2^{\circ}$, $s=1.8$, and $\rho=6$. The bottom row shows the results from a simulation of the data recorded at a longitude of $188.7^{\circ}$. The parameters used in this simulation were $\chi_{o}=-13^{\circ}$, $s=1$, and $\rho=3.8$. The contour levels shown in both rows are drawn at 0.2, 0.4, 0.6, and 0.8 of the joint density’s peak value.
@@
The top row of Figure 11 is a simulation of the PA-EA joint density at a pulse longitude of $178.1^{\circ}$ in PSR B1451-68. This simulation should be compared to the right center panel in Figure 2 of Dyks et al. (2021). The observed joint density consists of two circularly symmetric components. They reside at near-antipodal points close to the equatorial plane of the Poincaré sphere, indicating that the polarization modes are approximately orthogonal and predominantly linearly polarized. One component is more prominent than the other. The model parameters estimated from these observed properties and used in the simulation were $\chi_{o}=2^{\circ}$, $s=1.8$, and $\rho=6$. The results of the simulation are qualitatively consistent with what is observed. The simulation also resembles the PA-EA joint density observed in the trailing outrider of PSR B0329+54 (see the bottom right corner of Figure 2 in Edwards & Stappers 2004).
@@ 5 Comparison of the Statistical Properties of the Polarization Angles
A comparison of the statistical properties of the EA derived in Sections 2 and 3 with those of the PA underscores their similarities and differences. When the amplitude of the polarization vector is constant, the pdf of the PA is always symmetric about the intrinsic PA, $\psi_{o}$. Consequently, the mean of the pdf is always equal to $\psi_{o}$. The mean PA is not biased by the instrumental noise, provided the noise in the Stokes parameters $Q$ and $U$ is independent and equal in magnitude (Montier et al. 2015a, 2015b). The standard deviation of the pdf, $\sigma_{\psi}$, is independent of $\psi_{o}$ and varies inversely with the SNR in linear polarization. The pdf becomes Gaussian as the SNR becomes large, at which point the confidence limits of the pdf are well approximated by $\sigma_{\psi}$. In contrast, the pdf of the EA is generally asymmetric. The standard deviation of the pdf can depend upon the value of the intrinsic EA, $\chi_{o}$, and varies inversely with the SNR in total polarization. The measured EA is biased by the instrumental noise, particularly at low SNR and large values of $\lvert\chi_{o}\rvert$. Since the pdf is asymmetric, its confidence limits are also asymmetric about the mean EA. As the SNR increases, the pdf becomes Gaussian when the SNR and intrinsic EA satisfy the condition $s>3/\cos(2\chi_{o})$, at which point the mean EA is equal to $\chi_{o}$, $\sigma_{\chi}$ is independent of $\chi_{o}$, and the confidence limits of the pdf are well approximated by $\sigma_{\chi}$. The pdf is never symmetric when $\chi_{o}=\pi/4$.
The pdf of the PA is generally bimodal when the amplitude of the polarization vector fluctuates due to the superposition of incoherent OPMs. The two components of the pdf are separated precisely by $\Delta\psi=\pi/2$. Both components are symmetric about their respective centroids, with widths scaling inversely as the ratio of the polarization fluctuations relative to the instrumental noise, $\rho$. The pdf of the EA is also bimodal. The two components of the pdf are generally asymmetric about their respective peaks, and their separation tends to be less than $2\chi_{o}$, particularly at large values of $\lvert\chi_{o}\rvert$. As $\rho$ increases, the width of the components decreases, and their separation approaches $2\chi_{o}$. The pdf becomes unimodal when the polarization fluctuations are small relative to either the mean polarization or the instrumental noise.
@@ 6 SUMMARY
The Stokes parameters are covariant when the radio emission from pulsars is comprised of superposed incoherent OPMs. In this instance, the joint probability density of the polarization angles and the individual pdfs of the EA and PA are unimodal if the mean polarization exceeds the polarization fluctuations and are bimodal if the fluctuations exceed the mean. When the mean intensities of the modes are equal, the widths of the two components in the EA’s bimodal pdf vary inversely with the polarization fluctuations relative to the instrumental noise. The component separation approaches twice the intrinsic EA as the fluctuations increase.
Appendix A Confidence Limits of the EA